I took a quiz and i was wondering if i got this problem right, i did it as follows:
3x-2y=4
-3x -3x
-2y=4-3x
I left the problem like this because the teacher was asking us to isolate "y" so it fine if i left it as -2y=4-3x

Answer by mananth(16946)

y is not isolated as yet
-2y=4-3x
divide by -2
-2y/-2= 4/-2 -3x/-2
y=-2 +3x/2